OpenStudy (anonymous):

write a quadratic equation with the given roots: 1+√11i and 1-√11i

OpenStudy (anonymous):

\(2\pm 2\sqrt{11i}=-b\pm \sqrt{b^2-4ac}\) if a = 1 so let b = -2 and solve for c.

OpenStudy (anonymous):

i dont know about that man

OpenStudy (anonymous):

i keep getting x^2+2x+13

OpenStudy (anonymous):

look a the quadratic formula and work backwards.

OpenStudy (anonymous):

i dont know how to work the formula backwards. i thought there was an alternative way

OpenStudy (anonymous):

anyone here to help me :{

OpenStudy (jdoe0001):

have you done other cases when the roots weren't a complex number?

OpenStudy (anonymous):

no i dont think so

OpenStudy (jdoe0001):

keep in mind the website right now is very lagged

OpenStudy (anonymous):

im getting the answer x^2+2x+13

OpenStudy (anonymous):

yeah i can tell lol its really bad

OpenStudy (jdoe0001):

if you had a quadratic, could you obtain its roots?

OpenStudy (anonymous):

idk :(

OpenStudy (jdoe0001):

well..... have you done quadratic factoring at all?

OpenStudy (anonymous):

yeah i got x^2+1x+1x+2. but i dont know what to do about the 11i squared

OpenStudy (jdoe0001):

ok... well ... gimme one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so i got the answer lol

OpenStudy (jdoe0001):

\(\bf x=1+\sqrt{11}\ i\implies x-1-\sqrt{11}\ i=0\implies ({\color{blue}{ x-1-\sqrt{11}\ i}})=0\\ \quad \\ x=1-\sqrt{11}\ i\implies x-1+\sqrt{11}\ i=0\implies ({\color{blue}{ x-1+\sqrt{11}\ i}})=0\qquad thus\\ \quad \\ ({\color{blue}{ x-1-\sqrt{11}\ i}})({\color{blue}{ x-1+\sqrt{11}\ i}})=0\\ \quad \\ ({\color{blue}{ x-1-\sqrt{11}\ i}})({\color{blue}{ x-1+\sqrt{11}\ i}})=\textit{original polynomial}\\ -------------------------------\\ recall \implies {\color{blue}{ (a-b)(a+b) = a^2-b^2}}\qquad thus\\ \quad \\ (x-1-\sqrt{11}\ i)(x-1+\sqrt{11}\ i)=0\\ \quad \\ \implies [{\color{blue}{ (x-1)}}-(\sqrt{11}\ i)][{\color{blue}{ (x-1)}}+(\sqrt{11}\ i)]=0\\ \quad \\ \implies {\color{blue}{ (x-1)}}^2-(\sqrt{11}\ i)^2=0\)

OpenStudy (anonymous):

again, it was just a math typo. thank you for your time

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

\[if~ a~and~b~are~roots~of~an~equation, then \left( x-a \right)\left( x-b \right)=0~is~the~required~equation.\] \[\left\{ x-\left( 1+\sqrt{11}\iota \right) \right\}\left\{ x-\left( 1-\sqrt{11}\iota \right) \right\}=0\] \[\left\{ \left( x-1 \right)-\sqrt{11}\iota \right\}\left\{ \left( x-1 \right)+\sqrt{11}\iota \right\}=0\] solve it.